OpenStudy (anonymous):

Juan invested his savings in two investment funds. The$8000 that he invested in Fund A returned a 1% profit. The amount that he invested in Fund B returned a 5% profit. How much did he invest in Fund B, if both funds together returned a 4% profit?

OpenStudy (anonymous):

@hero this is the new question they gave me

hero (hero):

a = Fund A = 8000 b = Fund B c = Total Amount invested Juan invested 8000 in fund a so: 8000 + b = c This is the profit equation. He gets a 1% profit return on the 8000 invested in fund a. He gets 5% for fund b, and together he receives 4% profit on the total investment: .01(8000) + .05b = .04(c)

hero (hero):

So you have to solve this system: 8000 + b = c .01(8000) + .05b = .04c Remember: \(1\% = \frac{1}{100} = .01\) \(5\% = \frac{5}{100} = .05\)

OpenStudy (anonymous):

I got b=-1599.2 is that correct

hero (hero):

Nothing will be negative. You can't invest negative money.

hero (hero):

Try to understand the equations first.

OpenStudy (anonymous):

6400.8

OpenStudy (anonymous):

when I solve for .01(8000) + .05b = .04c it gives me a negative number

hero (hero):

This is a systems of equations. You have to solve both equations together.

OpenStudy (anonymous):

im not sure what to do im really confused

hero (hero):

Have you ever solved a system of equations before?

OpenStudy (anonymous):

yes the way your telling me to do it Im getting a negative number

OpenStudy (anonymous):

I Solve for b first using .01(8000) + .05b = .04c then I sub B into 8000 + b = c correct

hero (hero):

I haven't told you to do anything yet. I just set up the system. You started doing your own thing.

hero (hero):

What is 1% of 8000?

hero (hero):

That's the same as asking you to simplify .01(8000)

OpenStudy (anonymous):

80

hero (hero):

Good. We were able to get that far....so now we have 8000 + b = c 80 + .05b = .04c

hero (hero):

Now, multiply both sides of the second equation by 100

OpenStudy (anonymous):

100(80+.05)=100(.04) correct?

OpenStudy (anonymous):

b=0.8c-1600 c=1.25b+2000

hero (hero):

What have you done now? I said to multiply both sides of the second equation by 100, which you did, but you removed the variables...WHY?

hero (hero):

100(80+.05)=100(.04) <---- not correct 100(80 + .05b) = 100(.04c) <--- is correct

OpenStudy (anonymous):

b=1/2000c-0.0006

hero (hero):

You're doing your own thing again...

hero (hero):

Either that or you don't know how to simplify the equation after multiplying both sides by 1000.

OpenStudy (anonymous):

I don't know how to

OpenStudy (anonymous):

I keep coming up with the wrong answer so I don't know what I am doing

hero (hero):

After multiplying both sides by 100 you get 100(80 + .05b) = 100(.04c) After simplification you get 8000 + 5b = 4c

hero (hero):

That is still the second equation

hero (hero):

If we put both equations back together again we have 8000 + b = c 8000 + 5b = 4c

hero (hero):

The next thing to do is use substitution since: c = 8000 + b Substitute 8000 + b in place of c in the second equation

hero (hero):

Can you do that?

OpenStudy (anonymous):

am i solving for b

hero (hero):

Go back to the original question to find out what you're solving for.

OpenStudy (anonymous):

24000

hero (hero):

Wow...

hero (hero):

You actually got it

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

thank you so much

hero (hero):

I see what you were trying to do earlier now. You're solving for b, however, you have to first figure out if you plan to use elimination or substitution to solve the system. If you use elimination, you have to make sure you eliminate variable c. If you use substitution, you have to make sure the final equation you solve is in terms of variable b.

OpenStudy (anonymous):

yes I was very confused on what to do but thank you @hero